OpenStudy (tornmino):

A formula for finding the area of a rectangle is A=l⋅w. If you know the area (A) and length (l) of a rectangle, which formula can you use to find the width (w)?

OpenStudy (dustin_whitelock):

w=A/l, easy.

OpenStudy (tornmino):

Mind showing me how to do it so in the future I got it?

OpenStudy (dustin_whitelock):

divide both sides by W

OpenStudy (dustin_whitelock):

Or better do this....

OpenStudy (tornmino):

Thanks, but also I got another one I need help on, A formula for finding the perimeter of a rectangle is P=2L+2W. If you know the perimeter (P) and width (W) of a rectangle, which formula can you use to find the length (L)?

OpenStudy (dustin_whitelock):

It's a simple algebraic work, which doesn't have any step.

OpenStudy (dustin_whitelock):

2w=p-2l => w=p/2 - l

OpenStudy (tornmino):

Alright, thanks,

OpenStudy (dustin_whitelock):

Aight, if you find it helpful, give me a medal and close this question! ;)

OpenStudy (tornmino):

Got it.